### v1

@@ 2.1 Preliminaries: geometric classification
Fourthly, to prove ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and ${{\rm T}}^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. Let $c_{i,j}^{k},c_{i,j,k}^{l}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of ${\mathbb{V}}$. If there is a pair of maps $(f,(a_{i}^{j}))$, where $f:\mathbb{C}^{*}\to I$ and $a_{i}^{j}:\mathbb{C}^{*}\to\mathbb{C}$ are such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of ${\mathbb{V}}$ for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{i,j}^{k}(t),c_{i,j,k}^{l}(t)$ of $(\mu\big{(}f(t)\big{)},\mu^{\prime}\big{(}f(t)\big{)})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim\limits_{t\to 0}c_{i,j}^{k}(t)=c_{i,j}^{k}$ and $\lim\limits_{t\to 0}c_{i,j,k}^{l}(t)=c_{i,j,k}^{l}$, then ${{\rm T}}(*)\to{{\rm T}}^{\prime}$. In this case $E_{1}(t),\dots,E_{n}(t)$ and $f(t)$ are called a parametrized basis and a parametrized index for ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, respectively. Fithly, to prove ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$, we may use an analogous of Remark 15 for parametric families that follows from Lemma 16, see [4].
Lemma 16.
@@
Consider the family of Lie-Yamaguti (Bol) algebras ${\rm T}(*)$ and the Lie-Yamaguti (Bol) algebra ${\rm T}^{\prime}$. Suppose ${\rm T}(*)\to{\rm T}^{\prime}$. Let ${\mathcal{Z}}$ be a Zariski closed in $\mathcal{T}_{n}$ that is stable by the action of the invertible upper (lower) triangular matrices. Then if there is a representation $(\mu(\alpha),\mu^{\prime}(\alpha))$ of ${\rm T}(\alpha)$ in ${\mathcal{Z}}$ for every $\alpha\in I$, then there is a representation $(\lambda,\lambda^{\prime})$ of ${\rm T}^{\prime}$ in ${\mathcal{Z}}$.
@@
Remark 17.
Let $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ represent two Lie-Yamaguti algebras. Suppose $(\lambda,0)\not\in\overline{O((\mu,0))}$, ((((resp. $(0,\lambda^{\prime})\not\in\overline{O((0,\mu^{\prime}))}$)))), then $(\lambda,\lambda^{\prime})\not\in\overline{O((\mu,\mu^{\prime}))}$. As we construct the classification of Lie-Yamaguti (Bol) algebras from a certain class of algebras with a single multiplication which remains unchanged, this remark becomes very useful.
2.2 The geometric classification of Lie-Yamaguti algebras
In this subsection, we determine all the irreducible components of the variety of four-dimensional nilpotent Lie-Yamaguti algebras.
Theorem 18.
The variety of four-dimensional nilpotent Lie-Yamaguti algebras has dimension $14$ and it has $3$ irreducible components defined by
@@ 2.2 The geometric classification of Lie-Yamaguti algebras
Proof.
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
@@ 2.3 The geometric classification of Bol algebras
Proof.
From the proof of Theorem 18, we obtain that there are degeneration from the algebras $\mathcal{L}_{19}^{\alpha}$ and $\mathcal{L}_{29}$ to the algebras $\mathcal{L}_{04},$ $\mathcal{L}_{05},$ $\mathcal{L}_{06},$ $\mathcal{L}_{07},$ $\mathcal{L}_{08},$ $\mathcal{L}_{09},$ $\mathcal{L}_{10},$ $\mathcal{L}_{11},$ $\mathcal{L}_{12},$ $\mathcal{L}_{13},$ $\mathcal{L}_{14},$ $\mathcal{L}_{15},$ $\mathcal{L}_{16}^{\alpha},$ $\mathcal{L}_{17}^{\alpha},$ $\mathcal{L}_{18}^{\alpha},$ $\mathcal{L}_{22},$ $\mathcal{L}_{23}^{\alpha},$ $\mathcal{L}_{24},$ $\mathcal{L}_{25},$ $\mathcal{L}_{26},$ $\mathcal{L}_{27},$ $\mathcal{L}_{28}.$
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
@@
$\mathfrak{B}_{06}^{\alpha}$$\not\rightarrow$ $\mathfrak{B}_{10}$ =$\mathcal{R}=\left\{\begin{array}[]{l}c_{1,2,1}^{3}=0,\ c_{2,3,2}^{4}=0.\end{array}\right\}$
Here $c_{i,j}^{k}$ and $c_{i,j,k}^{l}$ coefficients are structural constants in the basis. ∎
@@ 2.4 The geometric classification of compatible Lie algebras
Proposition 20.
The variety of three-dimensional nilpotent compatible Lie algebras is irreducible and defined by $\mathcal{C}=\overline{O(\mathfrak{L}_{2}^{\alpha})},$ it has dimension $4$.
Proof.
@@
Theorem 21.
The variety of four-dimensional nilpotent compatible Lie algebras is irreducible and defined by $\mathcal{C}=\overline{O({L_{10}^{\alpha})}},$ it has dimension $13$.
Proof.
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have $\dim O(L_{10}^{\alpha})=13.$

### v2

@@ 2.1 Preliminaries: geometric classification
Secondly, to show degenerations, let $\rm T$}} and $\rm T$$^{\prime}$ be two Lie–Yamaguti (Bol) algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${{\mathcal{T}}}_{n}$, respectively. Let $c_{i,j}^{k},c_{i,j,k}^{l}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of $\mathbb{V}$}. If there exist $n^{2}$ maps $a_{i}^{j}(t):\mathbb{C}^{*}\to\mathbb{C}$ such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of $\mathbb{V}$} for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{i,j}^{k}(t),c_{i,j,k}^{l}(t)$ of $(\mu,\mu^{\prime})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim\limits_{t\to 0}c_{i,j}^{k}(t)=c_{i,j}^{k}$ and $\lim\limits_{t\to 0}c_{i,j,k}^{l}(t)=c_{i,j,k}^{l}$, then ${{\rm T}}\to{{\rm T}}^{\prime}$. In this case, $E_{1}(t),\dots,E_{n}(t)$ is called a parametric basis for ${{\rm T}}\to{{\rm T}}^{\prime}$. If $E_{e_{1}}^{t},E_{e_{2}}^{t},E_{e_{3}}^{t},E_{e_{4}}^{t}$ is a parametric basis for ${\bf A}\to{\bf B}$, then we denote a degeneration by ${\bf A}\xrightarrow{(E_{e_{1}}^{t},E_{e_{2}}^{t},E_{e_{3}}^{t},E_{e_{4}}^{t})}{\bf B}$.
Thirdly, to prove non-degenerations we may use a remark that follows from the following lemma (see [4] and the references therein).
Lemma 2.3.
@@
Consider two Lie–Yamaguti (Bol) algebras $\rm T$} and $\rm T$$^{\prime}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Let $\mathcal{Z}$} be a Zariski closed set in ${\mathcal{T}}_{n}$ that is stable by the action of the invertible upper (lower) triangular matrices. Then if there is a representation $(\mu,\mu^{\prime})$ of $\rm T$} in $\mathcal{Z}$}, then there is a representation $(\lambda,\lambda^{\prime})$ of $\rm T$$^{\prime}$ in $\mathcal{Z}$}.
To apply this lemma, we will give the explicit definition of the appropriate stable Zariski closed $\mathcal{Z}$} in terms of the variables $c_{i,j}^{k},c_{i,j,k}^{l}$ in each case. For clarity, we assume by convention that $c_{i,j}^{k}=0$ (resp. $c_{i,j,k}^{l}=0$) if $c_{i,j}^{k}$ (resp. $c_{i,j,k}^{l}$) is not explicitly mentioned in the definition of $\mathcal{Z}$}.
Remark 2.4.
Let $\rm T$}} and $\rm T$$^{\prime}$ be two Lie–Yamaguti (Bol) algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${{\rm T}}_{n}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Then if $\mu,\lambda$ represent algebras ${\rm T}_{0},{\rm T}^{\prime}_{0}$ in the affine space $\mathbb{C}^{n^{3}}$ and $\mu^{\prime},\lambda^{\prime}$ represent algebras ${\rm T}_{1},{\rm T}^{\prime}_{1}$ in the affine space $\mathbb{C}^{n^{4}}$ of algebras with a single multiplication, respectively, we have ${\rm T}_{0}\to{\rm T}^{\prime}_{0}$ and ${\rm T}_{1}\to{\rm T}^{\prime}_{1}$. So, for example, $(0,\mu)$ can not degenerate to $(\lambda,0)$ unless $\lambda=0$.
Fourthly, to prove ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and $\rm T$$^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. Let $c_{i,j}^{k},c_{i,j,k}^{l}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of $\mathbb{V}$}. If there is a pair of maps $(f,(a_{i}^{j}))$, where $f:\mathbb{C}^{*}\to I$ and $a_{i}^{j}:\mathbb{C}^{*}\to\mathbb{C}$ are such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of $\mathbb{V}$} for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{i,j}^{k}(t),c_{i,j,k}^{l}(t)$ of $(\mu\big{(}f(t)\big{)},\mu^{\prime}\big{(}f(t)\big{)})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim\limits_{t\to 0}c_{i,j}^{k}(t)=c_{i,j}^{k}$ and $\lim\limits_{t\to 0}c_{i,j,k}^{l}(t)=c_{i,j,k}^{l}$, then ${{\rm T}}(*)\to{{\rm T}}^{\prime}$. In this case $E_{1}(t),\dots,E_{n}(t)$ and $f(t)$ are called a parametrized basis and a parametrized index for ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, respectively. Fithly, to prove ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$, we may use an analogous of Remark 2.4 for parametric families that follows from Lemma 2.5, see [4].
Lemma 2.5.
Consider the family of Lie–Yamaguti (Bol) algebras ${\rm T}(*)$ and the Lie–Yamaguti (Bol) algebra $\rm T$$^{\prime}$. Suppose ${\rm T}(*)\to{\rm T}^{\prime}$. Let $\mathcal{Z}$} be a Zariski closed set in $\mathcal{T}_{n}$ that is stable by the action of the invertible upper (lower) triangular matrices. Then if there is a representation $(\mu(\alpha),\mu^{\prime}(\alpha))$ of ${\rm T}(\alpha)$ in $\mathcal{Z}$} for every $\alpha\in I$, then there is a representation $(\lambda,\lambda^{\prime})$ of $\rm T$$^{\prime}$ in $\mathcal{Z}$}.
@@
Remark 2.6.
Let $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ represent two Lie–Yamaguti algebras. Suppose $(\lambda,0)\not\in\overline{O((\mu,0))}$, ((resp. $(0,\lambda^{\prime})\not\in\overline{O((0,\mu^{\prime}))}$)), then $(\lambda,\lambda^{\prime})\not\in\overline{O((\mu,\mu^{\prime}))}$. As we construct the classification of Lie–Yamaguti (Bol) algebras from a certain class of algebras with a single multiplication which remains unchanged, this remark becomes very useful.
2.2 The geometric classification of Lie–Yamaguti algebras
In this subsection, we determine all the irreducible components of the variety of four-dimensional nilpotent Lie–Yamaguti algebras.
Theorem 2.7.
The variety of four-dimensional nilpotent Lie–Yamaguti algebras has dimension $14$ and it has $3$ irreducible components defined by
@@ 2.2 The geometric classification of Lie–Yamaguti algebras
Proof 2.8.
After carefully checking the dimensions of orbit closures of the more important algebras, we have
@@
As the orbit dimensions of these three algebras coincide, no degeneration occurs among them.
@@ 2.3 The geometric classification of Bol algebras
Proof 2.10.
From the proof of Theorem 2.7, we obtain that there are degenerations from the algebras $\mathcal{L}_{19}^{\alpha}$ and $\mathcal{L}_{29}$ to the algebras $\mathcal{L}_{04},$ $\mathcal{L}_{05},$ $\mathcal{L}_{06},$ $\mathcal{L}_{07},$ $\mathcal{L}_{08},$ $\mathcal{L}_{09},$ $\mathcal{L}_{10},$ $\mathcal{L}_{11},$ $\mathcal{L}_{12},$ $\mathcal{L}_{13},$ $\mathcal{L}_{14},$ $\mathcal{L}_{15},$ $\mathcal{L}_{16}^{\alpha},$ $\mathcal{L}_{17}^{\alpha},$ $\mathcal{L}_{18}^{\alpha},$ $\mathcal{L}_{22},$ $\mathcal{L}_{23}^{\alpha},$ $\mathcal{L}_{24},$ $\mathcal{L}_{25},$ $\mathcal{L}_{26},$ $\mathcal{L}_{27},$ $\mathcal{L}_{28}.$
After carefully checking the dimensions of orbit closures of the more important algebras, we have
@@
$\mathfrak{B}_{06}^{\alpha}$ $\not\rightarrow$ $\mathfrak{B}_{10}$ $\mathcal{R}=\left\{c_{1,2,1}^{3}=0,\ c_{2,3,2}^{4}=0\right\}$
Here, the coefficients $c_{i,j}^{k}$ and $c_{i,j,k}^{l}$ are structural constants in the basis $\{e_{1},e_{2},e_{3},e_{4}\}$.
@@ 2.4 The geometric classification of compatible Lie algebras
Proposition 2.11.
The variety of three-dimensional nilpotent compatible Lie algebras is irreducible and defined by $\mathcal{C}=\overline{O(\mathfrak{L}_{2}^{\alpha})}$, and the dimension of this variety is equal to $4$.
Proof 2.12.
@@
Theorem 2.13.
The variety of four-dimensional nilpotent compatible Lie algebras is irreducible and defined by $\mathcal{C}=\overline{O({L_{10}^{\alpha})}}$, and the dimension of this variety is equal to $13$.
Proof 2.14.
After carefully checking the dimensions of orbit closures of the more important algebras, we have $\dim O(L_{10}^{\alpha})=13.$